# An operator splitting analysis of Wasserstein–Fisher–Rao gradient flows

### v1

@@ B.1 Moment ODEs and analytic solutions
We collect here the derivation and exact solutions of the moment ODEs for WFR and sequential splitting. Before stating our results we recall the following properties of matrix exponentials: for $X,Y$ matrices we have
$e^{X}e^{Y}$ $=e^{X+Y}\quad\text{if}\quad XY=YX$
$e^{YXY^{-1}}$ $=Ye^{X}Y^{-1}\quad Y\enskip\text{invertible}$
$Xe^{X}$ $=e^{X}X.$
We first refine the result of [liero2025evolution, Remark 2.3] and obtain an explicit expression for the covariance matrix of the WFR flow rather than its inverse. The specific form derived below will be particularly useful to obtain the convergence rate of $C_{t}$ to $C_{\pi}$ in Lemma 3.1. Additionally, we obtain explicit solutions for the mean, which are not detailed in [liero2025evolution] and seem to be missing from the literature.
Throughout this section, we focus on PDEs of the form
$\partial_{t}\mu_{t}$ $=f_{\textrm{W}}(\mu_{t})+f_{\textrm{FR}}(\mu_{t})+\mu_{t}\left(g_{M}(\mu_{t})-\mathbb{E}_{\mu_{t}}[g_{M}]\right)$ (B.1)
$g_{M}(\mu_{t})$ $=\nabla g(\mu_{t})^{\top}M_{t}\nabla g(\mu_{t})$ (B.2)
initialised at $\mu_{0}=\mathcal{N}(x;m_{0},C_{0})$ with $g$ as defined in (2.4), $\pi=\mathcal{N}(x;m_{\pi},C_{\pi})$ and $M_{t}\in\mathbb{R}^{d\times d}$ a potentially time dependent symmetric matrix independent of $x,\mu_{t}$. The ODEs for the exact WFR given in [us, Appendix D] are obtained for $M_{t}\equiv 0$. Assume further that $C_{0}$ and $C_{\pi}$ are invertible (non-degenerate) covariance matrices. It holds straightforwardly that $\mu_{t}=\mathcal{N}(x;m_{t},C_{t})$ and the following lemmas characterise the evolution of $m_{t},C_{t}$ with time.
@@
Moment ODEs, Multivariate Gaussian. Consider (B.1) with the aforementioned conditions. Then $m_{t},C_{t}$ satisfy for all $t>0$,
$\frac{dC_{t}}{dt}$ $=-C_{t}C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})C_{t}-\left(I-\tfrac{1}{2}C_{\pi}+2M_{t}\right)C_{\pi}^{-1}C_{t}-C_{t}C_{\pi}^{-1}(I-\tfrac{1}{2}C_{\pi}+2M_{t})+2(I+M_{t})$ (B.3)
$\frac{dm_{t}}{dt}$ $=-(C_{t}C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})+(I+2M_{t})C_{\pi}^{-1})(m_{t}-m_{\pi})$ (B.4)
@@
By straightforward integration and assuming interchange of derivatives and integrals,
$\frac{dC_{t}}{dt}=\int(x-m_{t})(x-m_{t})^{\top}\partial_{t}\mu_{t}(x)dx$ $=\int(x-m_{t})(x-m_{t})^{\top}(f_{\textrm{W}}(\mu_{t})+f_{\textrm{FR}}(\mu_{t}))dx$ (B.5)
@@
The first term on the rhs is obtained directly from the covariance ODE for the exact WFR, see [us, Appendix D] for a derivation,
$\int(x-m_{t})(x-m_{t})^{\top}(f_{\textrm{W}}(\mu_{t})+f_{\textrm{FR}}(\mu_{t}))dx=-C_{t}C_{\pi}^{-1}C_{t}+C_{t}-C_{\pi}^{-1}C_{t}-C_{t}C_{\pi}^{-1}+2I.$ (B.6)
@@
$g_{M}(\mu_{t})=(x-m_{t}+m_{t}-\tilde{m}_{t})^{\top}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(x-m_{t}+m_{t}-\tilde{m}_{t})$ (B.7)
@@
$\int(x-m_{t})(x-m_{t})^{\top}g_{M}(\mu_{t})\mu_{t}(x)dx=\int(x-m_{t})(x-m_{t})^{\top}(x-m_{t})^{\top}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(x-m_{t})\mu_{t}(x)dx$
$+2\int(x-m_{t})(x-m_{t})^{\top}(x-m_{t})^{\top}\mu_{t}(x)dx\,\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(m_{t}-\tilde{m}_{t})$
$+C_{t}(m_{t}-\tilde{m}_{t})^{\top}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(m_{t}-\tilde{m}_{t})$
$=2C_{t}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}C_{t}+C_{t}Tr[C_{t}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}]+C_{t}(m_{t}-\tilde{m}_{t})^{\top}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(m_{t}-\tilde{m}_{t})$
$=2(C_{t}C_{\pi}^{-1}M_{t}C_{\pi}^{-1}C_{t}-C_{t}C_{\pi}^{-1}M_{t}-M_{t}C_{\pi}^{-1}C_{t}+M_{t})+C_{t}Tr[C_{t}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}]$
$+C_{t}(m_{t}-\tilde{m}_{\gamma})^{\top}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(m_{t}-\tilde{m}_{t})$ (B.8)
where in the second last line we have used that for a random variable $X\in\mathbb{R}^{d\times 1}$, $X\sim N(0,\Sigma)$ and any constant matrix $A$, $\mathbb{E}[XX^{\top}X^{\top}AX]=\mathbb{E}[XX^{T}AXX^{T}]=\Sigma(A+A^{\top})\Sigma+\Sigma Tr[\Sigma A]$. Also,
$\mathbb{E}_{\mu_{t}}[g_{M}]\int(x-m_{t})(x-m_{t})^{\top}\mu_{t}(x)dx=C_{t}\left(Tr[C_{t}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}]+(m_{t}-\tilde{m}_{t})^{\top}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(m_{t}-\tilde{m}_{t})\right)$ (B.9)
Substituting (B.8) and (B.9) into (B.5) and combining with (B.6) yields the covariance ODE (B.3).
The mean ODE is obtained in a similar way
$\frac{dm_{t}}{dt}=\int x\partial_{t}\mu_{t}(x)dx$ $=\int x(f_{\textrm{W}}(\mu_{t})+f_{\textrm{FR}}(\mu_{t}))dx+\int x(g_{M}(\mu_{t})-\mathbb{E}_{\mu_{t}}[g_{M}])\mu_{t}(x)dx,$
where the first term on the rhs is obtained directly from the mean ODE for the exact WFR, see [us, Appendix D] for derivation,
$\int x(f_{\textrm{W}}(\mu_{t})+f_{\textrm{FR}}(\mu_{t}))dx=-(C_{t}+I)C_{\pi}^{-1}(m_{t}-m_{\pi})$ (B.10)
and for the remaining term, using (B.7),
$\int x(g_{M}(\mu_{t})-\mathbb{E}_{\mu_{t}}[g_{M}])\mu_{t}(x)dx$ $=\int(x-m_{t})(x-m_{t}+m_{t}-\tilde{m}_{t})^{\top}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(x-m_{t}+m_{t}-\tilde{m}_{t})\mu_{t}(x)dx$
$=2C_{t}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(m_{t}-\tilde{m}_{t})$
$=2C_{t}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(m_{t}-\tilde{C}_{t}C_{\pi}^{-1}m_{\pi}+\tilde{C}_{t}C_{t}^{-1}m_{t})$
$=2C_{t}\tilde{C}_{t}^{-1}M_{t}C_{\pi}^{-1}m_{t}-2C_{t}\tilde{C}_{t}^{-1}M_{t}C_{\pi}^{-1}m_{\pi}$
$=2(C_{t}C_{\pi}^{-1}-I)M_{t}C_{\pi}^{-1}(m_{t}-m_{\pi}).$
Combining with (B.10) yields (B.4).
@@
Analytic Solutions, Multivariate Gaussian. Given the conditions of Lemma B.1 and $\Gamma:=C_{\pi}^{-1}+\tfrac{1}{2}I$. Then the mean and covariance ODEs (B.4) and (B.3) with initial conditions $m_{0},C_{0}$ have analytic solution for all $t>0$ given by
$C_{t}$ $=C_{\pi}+e^{-\Gamma t}E_{0}(I-(W_{t}^{-1}+E_{0})^{-1}E_{0})e^{-\Gamma t}$ (B.11)
$m_{t}$ $=m_{\pi}+(C_{t}-C_{\pi})e^{tC_{\pi}^{-1}}(C_{0}-C_{\pi})^{-1}(m_{0}-m_{\pi}).$ (B.12)
where $W_{t}:=(2I+C_{\pi})^{-1}-Z_{t}^{M}-e^{-\Gamma t}(2I+C_{\pi})^{-1}e^{-\Gamma t}$, $Z_{t}^{M}:=2\int_{0}^{t}e^{-\Gamma s}M_{s}e^{-\Gamma s}ds$ with $Z_{t}^{M}<\infty$ for all $t>0$ and $E_{0}:=C_{0}-C_{\pi}$. Notice in particular that $m_{t}$ is independent of $M_{t}$.
@@
$z_{t}=(C_{t}-C_{\pi})^{-1}(m_{t}-m_{\pi})$ (B.14)
and differentiating with respect to $t$ yields (using the shorthand notation $\frac{dC_{t}}{dt}=\dot{C}_{t}$,
$\dot{z_{t}}$ $=-(C_{t}-C_{\pi})^{-1}\dot{C}_{t}(C_{t}-C_{\pi})^{-1}(m_{t}-m_{\pi})+(C_{t}-C_{\pi})^{-1}\dot{m}_{t}$
$=-(C_{t}-C_{\pi})^{-1}\dot{C}_{t}z_{t}-(C_{t}-C_{t})^{-1}(C_{t}C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})+(I+2M_{t})C_{\pi}^{-1})(C_{t}-C_{\pi})z_{t}$
using (B.4) and $(C_{t}-C_{\pi})z_{t}=m_{t}-m_{\pi}$. Rearranging (B.3) as
$\dot{C}_{t}$ $=(2C_{t}C_{\pi}^{-1}M_{t}C_{\pi}^{-1}-2M_{t}C_{\pi}^{-1}-C_{t}C_{\pi}^{-1}-C_{\pi}^{-1})(C_{t}-C_{\pi})-(C_{t}-C_{\pi})C_{\pi}^{-1}$
$=(C_{t}C_{\pi}^{-1}(2M_{t}C_{\pi}^{-1}-I)-(I+2M_{t})C_{\pi}^{-1})(C_{t}-C_{\pi})-(C_{t}-C_{\pi})C_{\pi}^{-1}$
and substituting into the ODE for $z_{t}$ yields
$\dot{z_{t}}$ $=-(C_{t}-C_{\pi})^{-1}[C_{t}C_{\pi}^{-1}(2M_{t}C_{\pi}^{-1}-I)-(I+2M_{t})C_{\pi}^{-1}](C_{t}-C_{\pi})z_{t}+(C_{t}-C_{\pi})^{-1}(C_{t}-C_{\pi})C_{\pi}^{-1}z_{t}$
$-(C_{t}-C_{t})^{-1}(C_{t}C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})+(I+2M_{t})C_{\pi}^{-1})(C_{t}-C_{\pi})z_{t}$
$=C_{\pi}^{-1}z_{t}$
@@
which has analytic solution $z_{t}=e^{tC_{\pi}^{-1}}z_{0}$. Inserting (B.14) and rearranging yields (B.12).
For the covariance ODE, we first consider the so-called natural transformation as in [liero2025evolution] to obtain an analytic solution for $C_{t}^{-1}$. Using the transformation $A_{t}=C_{t}^{-1}$ and differentiating with respect to time yields
$\dot{A}_{t}$ $=-A_{t}\dot{C}_{t}A_{t}$
$=-C_{t}^{-1}\left(-C_{t}C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})C_{t}-\left(I-\tfrac{1}{2}C_{\pi}+2M_{t}\right)C_{\pi}^{-1}C_{t}-C_{t}C_{\pi}^{-1}(I-\tfrac{1}{2}C_{\pi}+2M_{t})+2(I+M_{t})\right)C_{t}^{-1}$
$=C_{\pi}^{-1}-2C_{\pi}^{-1}M_{t}C_{\pi}^{-1}+C_{t}^{-1}\left(I+2M_{t}\right)C_{\pi}^{-1}+C_{\pi}^{-1}(I+2M_{t})C_{t}^{-1}-C_{t}^{-1}-2C_{t}^{-1}(I+M_{t})C_{t}^{-1}$
$=C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})+A_{t}\left(I+2M_{t}\right)C_{\pi}^{-1}+C_{\pi}^{-1}(I+2M_{t})A_{t}-A_{t}-2A_{t}(I+M_{t})A_{t}$
Let $\Sigma_{t}:=A_{t}-C_{\pi}^{-1}$ and $\Gamma:=C_{\pi}^{-1}+\frac{1}{2}I$ and $B_{t}:=e^{\Gamma t}\Sigma_{t}e^{\Gamma t}$, then
$\dot{B}_{t}=e^{\Gamma t}(\Gamma\Sigma_{t}+\dot{\Sigma}_{t}+\Sigma_{t}\Gamma)e^{\Gamma t}$ (B.15)
and
$\Gamma\Sigma_{t}+\dot{\Sigma}_{t}+\Sigma_{t}\Gamma$ $=\left(C_{\pi}^{-1}+\tfrac{1}{2}I\right)(A_{t}-C_{\pi}^{-1})+(A_{t}-C_{\pi}^{-1})\left(C_{\pi}^{-1}+\tfrac{1}{2}I\right)$
$+C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})+A_{t}\left(I+2M_{t}\right)C_{\pi}^{-1}+C_{\pi}^{-1}(I+2M_{t})A_{t}-A_{t}-2A_{t}(I+M_{t})A_{t}$
$=C_{\pi}^{-1}A_{t}+A_{t}C_{\pi}^{-1}-2C_{\pi}^{-2}-C_{\pi}^{-1}+A_{t}+A_{t}\left(I+2M_{t}\right)C_{\pi}^{-1}+C_{\pi}^{-1}(I+2M_{t})A_{t}$
$+C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})-A_{t}-2A_{t}(I+M_{t})A_{t}$
$=-2\Sigma_{t}(I+M_{t})\Sigma_{t}$
and inserting into (B.15) yields
$\dot{B}_{t}$ $=-2e^{\Gamma t}\Sigma_{t}(I+M_{t})\Sigma_{t}e^{\Gamma t}=-2B_{t}e^{-\Gamma t}(I+M_{t})e^{-\Gamma t}B_{t}$
Consider a further transformation, $D_{t}=B_{t}^{-1}$, then
$\dot{D}_{t}=-D_{t}\dot{B}_{t}D_{t}=2e^{-\Gamma t}(I+M_{t})e^{-\Gamma t}$ (B.16)
whose analytic solution has the representation
$D_{t}=D_{0}+2\int_{0}^{t}e^{-\Gamma s}(I+M_{s})e^{-\Gamma s}ds=(C_{0}^{-1}-C_{\pi}^{-1})^{-1}+Z_{t}$
with $Z_{t}:=2\int_{0}^{t}e^{-\Gamma s}(I+M_{s})e^{-\Gamma s}ds$ and since $D_{0}=(C_{0}^{-1}-C_{\pi}^{-1})^{-1}$. Finally,
$C_{t}^{-1}=C_{\pi}^{-1}+\Sigma_{t}=C_{\pi}^{-1}+e^{-\Gamma t}D_{t}^{-1}e^{-\Gamma t}=C_{\pi}^{-1}+e^{-\Gamma t}((C_{0}^{-1}-C_{\pi}^{-1})^{-1}+Z_{t})^{-1}e^{-\Gamma t}.$ (B.17)
Note that for the special case $M_{t}=0$, $Z_{t}=2\int_{0}^{t}e^{-2\Gamma s}ds=\Gamma^{-1}(I-e^{-2\Gamma t})$ so that the above coincides with the expression for $C_{t}^{-1}$ in [liero2025evolution, Remark 2.3] with $\alpha=\beta=1$ there.

### v2

@@ B.1 Moment ODEs and analytic solutions
We first refine the result of [LMT+25, Remark 2.3] and obtain an explicit expression for the covariance matrix of the WFR flow rather than its inverse. The specific form derived below will be particularly useful to obtain the convergence rate of $C_{t}$ to $C_{\pi}$ in Lemma 3.1. Additionally, we obtain explicit solutions for the mean, which are not detailed in [LMT+25] and seem to be missing from the literature. Throughout this section, we focus on PDEs of the form
$\partial_{t}\mu_{t}$ $=f_{\textrm{W}}(\mu_{t})+f_{\textrm{FR}}(\mu_{t})+\mu_{t}\left(g_{M}(\mu_{t})-\mathbb{E}_{\mu_{t}}[g_{M}]\right)$ (2.1)
$g_{M}(\mu_{t})$ $=\nabla g(\mu_{t})^{\top}M_{t}\nabla g(\mu_{t})$ (2.2)
initialised at $\mu_{0}$ with $g$ as defined in (2.4) and $M_{t}\in\mathbb{R}^{d\times d}$ a potentially time dependent symmetric matrix independent of $x,\mu_{t}$. The ODEs for the exact WFR given in [CP25, Appendix D] are obtained for $M_{t}\equiv 0$. Assume further that $C_{0}$ and $C_{\pi}$ are invertible (non-degenerate) covariance matrices. It holds straightforwardly that $\mu_{t}=\mathcal{N}(x;m_{t},C_{t})$ and the following lemmas characterise the evolution of $m_{t},C_{t}$ with time.
@@
Moment ODEs, Multivariate Gaussian. Consider (2.1) with the aforementioned conditions. Then $m_{t},C_{t}$ satisfy for all $t>0$,
$\frac{dC_{t}}{dt}$ $=-C_{t}C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})C_{t}-\left(I-\tfrac{1}{2}C_{\pi}+2M_{t}\right)C_{\pi}^{-1}C_{t}-C_{t}C_{\pi}^{-1}(I-\tfrac{1}{2}C_{\pi}+2M_{t})+2(I+M_{t})$ (2.3)
$\frac{dm_{t}}{dt}$ $=-(C_{t}C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})+(I+2M_{t})C_{\pi}^{-1})(m_{t}-m_{\pi})$ (2.4)
@@
By integration and assuming interchange of derivatives and integrals,
$\frac{dC_{t}}{dt}=\int(x-m_{t})(x-m_{t})^{\top}\partial_{t}\mu_{t}(x)dx$ $=\int(x-m_{t})(x-m_{t})^{\top}(f_{\textrm{W}}(\mu_{t})+f_{\textrm{FR}}(\mu_{t}))dx$ (2.5)
@@
The first term on the rhs is obtained directly from the covariance ODE for the exact WFR, see [CP25, Appendix D] for a derivation,
$\int(x-m_{t})(x-m_{t})^{\top}(f_{\textrm{W}}(\mu_{t})+f_{\textrm{FR}}(\mu_{t}))dx=-C_{t}C_{\pi}^{-1}C_{t}+C_{t}-C_{\pi}^{-1}C_{t}-C_{t}C_{\pi}^{-1}+2I.$ (2.6)
@@
$g_{M}(\mu_{t})=(x-m_{t}+m_{t}-\tilde{m}_{t})^{\top}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(x-m_{t}+m_{t}-\tilde{m}_{t})$ (2.7)
@@
$\int(x-m_{t})(x-m_{t})^{\top}g_{M}(\mu_{t})\mu_{t}(x)dx=2(C_{t}C_{\pi}^{-1}M_{t}C_{\pi}^{-1}C_{t}-C_{t}C_{\pi}^{-1}M_{t}-M_{t}C_{\pi}^{-1}C_{t}+M_{t})$
$+C_{t}Tr[C_{t}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}]+C_{t}(m_{t}-\tilde{m}_{\gamma})^{\top}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(m_{t}-\tilde{m}_{t})$ (2.8)
using that for a random variable $X\in\mathbb{R}^{d\times 1}$, $X\sim N(0,\Sigma)$ and any constant matrix $A$, $\mathbb{E}[XX^{\top}X^{\top}AX]=\mathbb{E}[XX^{T}AXX^{T}]=\Sigma(A+A^{\top})\Sigma+\Sigma Tr[\Sigma A]$. Also,
$\mathbb{E}_{\mu_{t}}[g_{M}]\int(x-m_{t})(x-m_{t})^{\top}\mu_{t}(x)dx=C_{t}\left(Tr[C_{t}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}]+(m_{t}-\tilde{m}_{t})^{\top}\tilde{C}_{t}^{-1}M_{t}\tilde{C}_{t}^{-1}(m_{t}-\tilde{m}_{t})\right)$ (2.9)
Substituting (2.8) and (2.9) into (2.5) and combining with (2.6) yields the covariance ODE (2.3). The mean ODE is obtained in a similar way and its proof is therefore omitted.
@@
Analytic Solutions, Multivariate Gaussian. Given the conditions of Lemma B.1 and $\Gamma:=C_{\pi}^{-1}+\tfrac{1}{2}I$. Then the mean and covariance ODEs (2.4) and (2.3) with initial conditions $m_{0},C_{0}$ have analytic solution for all $t>0$ given by
$C_{t}$ $=C_{\pi}+e^{-\Gamma t}E_{0}(I-(W_{t}^{-1}+E_{0})^{-1}E_{0})e^{-\Gamma t}$ (2.10)
$m_{t}$ $=m_{\pi}+(C_{t}-C_{\pi})e^{tC_{\pi}^{-1}}(C_{0}-C_{\pi})^{-1}(m_{0}-m_{\pi}).$ (2.11)
where $W_{t}:=(2I+C_{\pi})^{-1}-Z_{t}^{M}-e^{-\Gamma t}(2I+C_{\pi})^{-1}e^{-\Gamma t}$, $Z_{t}^{M}:=2\int_{0}^{t}e^{-\Gamma s}M_{s}e^{-\Gamma s}ds$ with $Z_{t}^{M}<\infty$ for all $t>0$ and $E_{0}:=C_{0}-C_{\pi}$. Notice that $m_{t}$ is independent of $M_{t}$.
@@
$z_{t}=(C_{t}-C_{\pi})^{-1}(m_{t}-m_{\pi})$ (2.13)
and differentiating with respect to $t$ and using (2.4) and $(C_{t}-C_{\pi})z_{t}=m_{t}-m_{\pi}$ yields
$\dot{z_{t}}$ $=-(C_{t}-C_{\pi})^{-1}\dot{C}_{t}z_{t}-(C_{t}-C_{t})^{-1}(C_{t}C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})+(I+2M_{t})C_{\pi}^{-1})(C_{t}-C_{\pi})z_{t}$
Rearranging (2.3) as
$\dot{C}_{t}$ $=(C_{t}C_{\pi}^{-1}(2M_{t}C_{\pi}^{-1}-I)-(I+2M_{t})C_{\pi}^{-1})(C_{t}-C_{\pi})-(C_{t}-C_{\pi})C_{\pi}^{-1}$
@@
and substituting into the ODE for $z_{t}$ yields $\dot{z}_{t}=C_{\pi}^{-1}z_{t}$, which has analytic solution $z_{t}=e^{tC_{\pi}^{-1}}z_{0}$. Inserting (2.13) and rearranging yields (2.11).
For the covariance ODE, consider the transformation used in [LMT+25, Remark 2.3],
$A_{t}=C_{t}^{-1}=C_{\pi}^{-1}+e^{-\Gamma t}B_{t}e^{-\Gamma t}$ (2.14)
Differentiating and rearranging yields
$\dot{A}_{t}$ $=C_{\pi}^{-1}(I-2M_{t}C_{\pi}^{-1})+A_{t}\left(I+2M_{t}\right)C_{\pi}^{-1}+C_{\pi}^{-1}(I+2M_{t})A_{t}-A_{t}-2A_{t}(I+M_{t})A_{t}$
$\dot{B}_{t}$ $=-2B_{t}e^{-\Gamma t}(I+M_{t})e^{-\Gamma t}B_{t}$
Again as in [LMT+25], setting $D_{t}=B_{t}^{-1}$ and differentiating with respect to $t$ yields an ODE whose exact solution is given by
$D_{t}=D_{0}+2\int_{0}^{t}e^{-\Gamma s}(I+M_{s})e^{-\Gamma s}ds=(C_{0}^{-1}-C_{\pi}^{-1})^{-1}+Z_{t},$